a fund drive announces that it has raised 5075, which is 29% of its goal. How much more must be raised?
Answer by mananth(16949)

let the amount be x
29 % is 5075
0.29x=5705
x= 5705/0.29
x= 19,672.4
Balance = 19,672-5075 =14,597